Debbie McAdams paid 4% interest on a $6,500 loan balance. Jan Burke paid $7,130 interest on a $57,500 loan. Based on 1 year:

a. What was the amount of interest paid by Debbie? $260



b. What was the interest rate paid by Jan? (Round your answer to the nearest tenth percent.) 12.4%



c. Debbie and Jan are both in the 28% tax bracket. Since the interest is deductible, how much would Debbie and Jan each save in taxes? (Round your answers to the nearest cent.)

I need help figuring this out and there is no sample in my book.
1
Expert's answer
2020-03-19T08:46:05-0400

a) amount of interest paid by Debbie = 4% of $6,500"=\dfrac{4}{100}\times \$6500=\$260"


b)interest rate paid by Jan="\frac{interest \ amount }{loan\ amount }\times100 =\frac{\$7130}{\$57500}\times100 \%=12.4
\%"


c) tax saved by Debbie = 28% of interest paid = "\frac{28}{100}\times \$260=\$72.8"


tax saved by Jan =28% of interest paid = "\frac{28}{100}\times \$7130=\$1996.4"